OpenStudy (anonymous):

If y varies inversely as x and the constant of variation is 30, what is y when x = 6 ?

OpenStudy (ash2326):

y varies inversely to x y=k/x k=constant=30 (given) y=30/x when x=6 y=30/6 y=5

OpenStudy (king):

5

OpenStudy (anonymous):

cinco...5